1. Let r(t) = -i-e2t j + (t? + 2t)k be the position of a particle moving in space. a. Find the particles velocity, speed andwhat is the answer for number 4
Answer #1

f(x,y,z)=In (x²+y? +z–4) f(0,-1,2)=ln (oP+(-1) +22–4]=In 1=0 |(2,3,-1)=ln (22+(3)2+(-1)2–4]=In 10 (6) Domain off Function f i